### v2

# Surprisingly High Redundancy in Electronic Structure Data

Abstract
Accurate prediction of electronic structure underpins advances in chemistry, materials science, and condensed matter physics. In recent years, Machine Learning (ML) has enabled the development of powerful surrogate models that can enable the prediction of the ground state electron density and related properties at a fraction of the computational cost of conventional first principles simulations. Such ML models typically rely on massive datasets generated through expensive Kohn-Sham Density Functional Theory calculations. A key reason for relying on such large datasets is the lack of prior knowledge about which portions of the data are essential, and which are redundant. This study reveals significant redundancies in electronic structure datasets across various material systems, including molecules, simple metals, and chemically complex alloys — challenging the notion that extensive datasets are essential for accurate ML-based electronic structure predictions. We demonstrate that even random pruning can substantially reduce dataset size with minimal loss in predictive accuracy. Furthermore, a state-of-the-art coverage-based pruning strategy that selects data across all learning difficulties, retains chemical accuracy and model generalizability using up to $100$-fold less data, while reducing training time by threefold or greater. By contrast, widely used importance-based pruning methods, which eliminate easy-to-learn data, can catastrophically fail at higher pruning factors due to significant reduction in data coverage. This heretofore unexplored high redundancy in electronic structure data holds the potential to identify a minimal, essential dataset representative of each material class.
I Introduction
The calculation and analysis of the electronic structure of materials underpins much of modern computational materials science. From predicting thermodynamic stability and charge transport to modeling chemical reactivity, the electronic structure serves as a fundamental description of how interacting atoms and electrons give rise to observable properties. Consequently, electronic structure calculations often serve as an important building block of multiscale materials modeling [tadmor2011modeling] with a significant share of contemporary high performance computing resources being devoted to such simulations [hafner2008ab, curtarolo2013high, lejaeghere2016reproducibility].
Kohn-Sham Density Functional Theory (KS-DFT) serves as the primary workhorse of modern electronic structure calculations [kohn1965self, hohenberg1964inhomogeneous, hafner2006toward]. Although formulated in terms of the so-called Kohn-Sham orbitals, the fundamental unknown in KS-DFT is the electron density. Indeed, many ground state material properties, e.g. structural parameters and elastic constants, may be computed from the electron density field, and it also serves as a starting point for ab initio descriptions of excited state phenomena. However, practical calculations using KS-DFT, especially for complex systems such as disordered alloys or materials with extended defects, can be limited by the cubic scaling computational cost of the method with respect to system size. As such, a considerable share of research has been devoted to the development of specialized solution techniques and high performance computing algorithms to overcome this bottleneck [gavini2023roadmap, goedecker1999linear, banerjee2016chebyshev, banerjee2018two, motamarri2014subquadratic, lin2014siesta, dogan2023real, banerjee2021ab, banerjee2016cyclic].
In recent years, Machine Learning (ML) has enabled the development of powerful surrogate models that can enable the prediction of the ground state electron density and related properties at a fraction of the computational cost of conventional first principles simulations [lewis2021learning, jorgensen2022equivariant, zepeda2021deep, chandrasekaran2019solving, fiedler2023predicting, brockherde2017bypassing, del2023deep, schleder2019dft, kulik2022roadmap, pathrudkar2022machine, pathrudkar2024electronic, pathrudkar2024HEA]. These ML models are typically trained on KS-DFT data and use descriptors of the atomic environment to produce the electron density as the output. Recent studies have demonstrated the success of such ML models in predicting electron densities across a wide range of systems, including pure metals, alloys, insulators, organic molecules and nanostructures  [zepeda2021deep, li2025image, chandrasekaran2019solving, pathrudkar2024electronic, pathrudkar2024HEA, pathrudkar2022machine]. Moreover, these models show promising extrapolation capabilities, accurately predicting electron densities for unseen compositions in complex alloys, atomic configurations with defects and dislocations, and for chemical species beyond training [li2025image, pathrudkar2024HEA, pathrudkar2024electronic]. The ML models have also been continuously improved to address several challenges, including the enforcement of physical symmetries, incorporation of uncertainty quantification capabilities, and improvements in training and prediction efficiency [jorgensen2022equivariant, teh2021machine, pathrudkar2024electronic].
The paradigm shift brought on by the development of ML based electronic structure techniques has motivated the creation of extensive datasets [ramakrishnan2014quantum, nandi2023multixc, pinheiro2020machine, liang2019qm] — sometimes comprising millions of structures — under the implicit assumption that large, exhaustive sampling is essential to achieve high predictive accuracy and generalizability [faber2017prediction, smith2018less, musil2019fast]. While this assumption has driven progress in materials informatics [jain2013commentary, saal2013materials, jain2020materials], it also raises questions about the efficiency and necessity of such data collection efforts, particularly given the computational cost and environmental footprint associated with high-throughput quantum mechanical calculations.
In this work, we examine the following fundamental but largely overlooked questions: How much of this electronic structure data is genuinely essential? Furthermore, how can such essential datasets be systematically obtained? To answer these questions, we methodically probe the redundancy in electron density data spanning molecular systems, elemental metals, and chemically complex concentrated alloys, and show that a surprisingly small subset is often sufficient to train ML models to within chemical accuracy. In the context of this work, we define redundancy as the presence of data points whose removal does not significantly impact ML model accuracy, generalizability, or derived physical quantities. Our findings challenge the prevailing view that predictive fidelity and generalizability require vast amounts of data and suggest that it is possible to identify minimal datasets that dramatically reduce computational cost without sacrificing accuracy. Thus our exploration of the redundancy of electronic structure data can potentially lead to a small foundational dataset for each material class, which in turn can be used for benchmarking and training foundation models.
The exploration of redundancy in chemical datasets has received prior attention in the literature [li2023exploiting, li2024md, chen2024beyond]. For instance Li et. al. [li2023exploiting] have demonstrated a high degree of redundancy in datasets of scalar material properties such as the formation energy and band gaps, and have proposed an iterative strategy to reduce the size of training data. However, there has not been a systematic exploration of redundancies in the electronic structure itself, particularly in the context of ML models. Notable prior efforts include the observation that electronic structure data for various simple systems often lie on low-dimensional manifolds, enabling reconstruction via low-rank tensors [blesgen2012approximation, motamarri2016tucker] or a small number of principal component modes [teh2021machine, pathrudkar2022machine]. Somewhat relatedly, recent contributions have also exploited redundancies in the electron density to develop strategic sampling schemes for grid-based representations [feng2025efficient, focassio2023linear], although the effectiveness of such approaches may depend sensitively on the manner in which core states are modeled. Thus, although previous studies have addressed redundancy in various contexts, statistically grounded one-shot pruning methods — free from ad hoc heuristics — have yet to be explored for electronic structure data. Accordingly, determining a statistically justifiable level of data redundancy remains an open problem, which we address here. Thus, the techniques explored here are applicable to any class of material without the need for domain knowledge and possibly even to fields beyond the electron density, e.g. the electronic wavefunctions or the local density of states.
The observation of redundancy in electronic structure data is closely connected to Kohn’s nearsightedness principle [prodan2005nearsightedness, kohn1996density]. Indeed, this result has been used to deduce that the single particle density matrix that appears in electronic structure calculations can be very sparse [benzi2007decay, benzi2013decay], and has led to a large number of electronic structure codes that exploit it [gavini2023roadmap, baer1997sparsity, suryanarayana2017nearsightedness, mohr2015accurate, shimojo2001linear, SURYANARAYANA2018288, nakata2020large, garcia2009linear, goedecker1999linear, skylaris2005introducing]. In the realm of ML-based electronic structure predictions, similar considerations can be used to motivate the prunability of electronic structure data: since local environments largely determine the electronic structure, many configurations differing only in distant parts of the system are expected to have nearly identical electronic descriptors. This implies substantial redundancy in datasets that exhaustively enumerate configurations, especially in large and complex systems. However, there has not been a systematic exploration of this aspect, particularly in the context of state-of-the-art neural network based ML models of electronic structure.
Neural network performance generally improves with data and model size, often following a power law scaling [kaplan2020scaling, hestness2017deep]. However, beyond a certain level, marginal accuracy gains require significantly more data [sorscher2022beyond, zhai2022scaling]. Thus, power law scaling suggests that much of the training data is redundant and can be pruned without notable performance loss [sorscher2022beyond].Commonly used approaches for selecting a smaller, representative subset of a large dataset that preserves ML model performance are referred to as data pruning techniques or coreset selection methods. These techniques are often broadly categorized into the following three major approaches [tan2024data], importance score based methods [paul2021deep, toneva2018empirical, tan2024data, he2024large], coverage driven methods [zheng2023_CCS, sener2018coreset, xia2022moderate] and optimization-based methods [yang2022dataset, mirzasoleiman2020coresets, killamsetty2021grad, killamsetty2021glister, killamsetty2021retrieve].
Most existing pruning methods have been evaluated primarily on classification problems, whereas electronic structure prediction is a regression task. Therefore, identifying the degree of prunability of electronic structure data, as well as the most suitable pruning technique remain unexplored. In the present work, two state-of-the-art methods: GraNd [paul2021deep] and Coverage-Centric Coreset Selection (CCS) [zheng2023_CCS] are adopted. Optimization-based methods are not considered in this work because, despite their strong theoretical foundation and impressive performance, they are computationally quite expensive, and can be challenging to implement on large-scale datasets [yang2024mind]. The GraNd technique is adopted because of its intuitive approach, ease of implementation, and ability to identify important examples early in the training. GraNd, an importance-score based method, assigns lower scores to easy-to-learn examples, which are deemed redundant and removed while pruning. This can lead to poor data coverage at high pruning factors, as only difficult-to-learn examples are retained, causing severe performance drop [zheng2023_CCS, guo2022deepcore]. CCS, a coverage-driven method, overcomes this limitation by proposing a probability density-based coverage strategy to improve the coverage of the coreset [zheng2023_CCS]. CCS is adopted here due to its exceptional performance at high pruning factors and its flexibility in incorporating any importance score.
Through the use of these data-pruning techniques, we explore the degree of redundancy in electronic structure data across a wide variety of material systems — a simple metal, chemically disordered concentrated complex alloy systems, and molecules. We show that high degrees of redundancy appear to be a generic feature of electronic structure data: even random pruning can substantially reduce dataset size with minimal loss in predictive accuracy, while state-of-the-art coverage-based methods unveil even more dramatic levels of redundancy. Thus, a remarkably large fraction of the dataset across all the materials studied can be systematically pruned while preserving prediction accuracy and model generalizability. We have further quantified the quality of predictions via calculations of the ground-state energy from predicted electron densities and have found that chemically accurate predictions (i.e., energy errors less than $1$ kcal/mol or $1.6$ mHa/atom) can be made even after pruning 9999% of the data. In particular, our results for complex disordered metallic alloys hold across composition space, thus suggesting the potential to identify a minimal, essential dataset representative of each material class. Additionally, we found that for the materials considered here, ML model training time on the pruned datasets can be reduced by a factor of three or more, thus enabling accelerated hyperparameter tuning and foundation model development.
The remainder of this paper is structured as follows: Section II presents our findings, followed by a discussion in Section III, and methodology details in Section IV.
II Results
In this section, we quantify the redundancy in electronic structure data for four different materials. We present the error in the ML model for various pruning factors to identify the fraction of data that can be pruned without significantly compromising the model’s performance and generalizability, and further quantify this through the accuracy of the energy computed from the ML-predicted electron density.
The four material systems considered here are pure aluminum, bulk water, and two different alloy systems across composition space — SiGeSn and CrFeCoNi. The first two serve as representative examples of simple metals and molecular systems, respectively, while the latter two illustrate technologically important chemically complex concentrated alloys. Notably our results for the alloy systems (discussed below) are particularly striking since the ML predictions are carried out over the entire composition space of each alloy [pathrudkar2024electronic].
The dataset consists of snapshots of KS-DFT simulation taken at intermittent time steps, with each snapshot capturing the atomic coordinates, species information, and the resulting electron density over the simulation cell. Given a grid point in the simulation domain, the ML model provides a map to the electron density at that point, from the nearby atomic configuration. Details of the KS-DFT simulations and the ML map are provided in Section IV.1.
II.1 Redundancy in diverse material datasets
Figure 1: Error in ML-predicted electron density for the original dataset, 90% and 99% randomly, and CCS-based pruned datasets. Each ML model was trained three times and the mean error is reported.
Figure 2: Error in energy with respect to KS-DFT, as obtained from the ML-predicted electron density for the original dataset, 90% CCS, and 99% CCS-based pruned datasets. The electron density prediction from one of the ML models was postprocessed in each case.
To explore the redundancy, pruning is performed by removing a subset of the grid points of the original dataset. We explore three approaches for data pruning: random grid-point-wise pruning, GraNd score-based pruning, and CCS-based pruning. In random grid-point-wise pruning, grid points from the original dataset are removed randomly. The importance score (i.e. GraNd score) is computed for each datapoint and used for the GraNd and CCS approaches. Upon training ML models on the pruned datasets, the errors in the predicted electron density on a testing dataset are computed. We report the commonly used normalized root mean squared error (NRMSE) [teh2021machine] as an error metric, defined in the supplemental materials (section V.1).
Figure 1 highlights the redundancy in electronic structure data. For all four systems, $90\%$ of the data can be pruned through the CCS-method, without compromising accuracy in the predicted electron density. Even $99\%$ redundancy can be identified by the CCS method: at such a high pruning factor, the errors increase as expected, but the predicted electron densities still result in chemically accurate energies (i.e., errors less than $1$ kcal/mole or $1.6$ mHa/atom) across all system, as shown in Figure 2. The total energy of the system is calculated from the ML-predicted electron density, via postprocessing [pathrudkar2024HEA]. Notably, while the errors in the electron density and the corresponding energy errors are correlated, the one-to-one relationship between them can break down when the density errors and hence the energy errors are themselves low. This situation can arise for low pruning percentages across all models, or for very high pruning percentages, for some models (e.g. those involving the alloy systems considered in this study). An additional issue to keep in mind is that in general, the CrFeCoNi alloy system is more difficult to predict accurately through ML models, because of the involvement of harder pseudopotentials, the inclusion of semi-core electrons for transition metals in the KS-DFT calculations, and the large amount of compositional variation [pathrudkar2024HEA]. Thus, the overall ML predictions are slightly less accurate for this system, as opposed to the other systems studied here. However, our key finding, in this regard is that there is still a high degree of redundancy in the CrFeCoNi electronic structure data, and the elimination of this redundancy does not reduce ML model prediction accuracy further.
Figure 3: (a) $H^{1}$ seminorm and (b) $H^{1}$ norm of the error for the ML-predicted electron density obtained from the original dataset, $90\%$, and $99\%$ randomly and CCS based pruned dataset.
Figure 4: Two-dimensional (2D) slices showing electron density obtained by KS-DFT, ML model trained on the original dataset, 90% CCS, and 99% CCS based pruned dataset for (a) Aluminum at 1500K (b) Water at 600K (c) SiGeSn at 2400K (d) CrFeCoNi for 5000K. The electron density($\rho$) along the solid line is also compared with the KS-DFT and the ML-predicted electron densities show remarkable agreement with the KS-DFT, true for all three ML models across four systems. The unit of electron density is $\text{Bohr}^{-3}$ in atomic units.
To demonstrate that the ML models not only capture the electron density field but also smooth spatial variations in it, we plot the $H^{1}$ seminorm and $H^{1}$ norm of the error in electron density field in Figure 3 (see the supplemental materials, section V.1 for further details). The errors in smoothness exhibit a trend similar to that of the electron density shown in Figure 1. For all four material systems, the $90\%$ CCS-pruned dataset achieves a similar degree of spatial smoothness as the original dataset (i.e., similar $H^{1}$ seminorm and $H^{1}$ norm values), whereas the $99\%$ CCS-pruned dataset shows a slight decrease. The $90\%$ and $99\%$ randomly pruned dataset has higher $H^{1}$ seminorm and $H^{1}$ norm values than CCS based pruning, for all four systems.
We compare the electron density fields obtained from KS-DFT with those predicted by ML models trained on three datasets — the original dataset, the $90\%$, and $99\%$ CCS-based pruned dataset — for all four material systems, as shown in Figure 4. ML models trained on the pruned datasets show excellent agreement with those trained on the original dataset and with KS-DFT calculations across all systems, confirming the high degree of lossless pruning. More details on this comparison are provided in the supplemental materials (section V.7).
II.2 Comparative assessment of pruning strategies
Figure 5: The error in the electron density prediction for various pruning factors for (a) Aluminum (b) Water (c) SiGeSn (d) CrFeCoNi. The shaded region represents the range (maximum to minimum) of three ML models around the mean, shown by a solid line.
Figure 6: The error in the total energy obtained via postprocessing from the predicted electron density for various pruning factors for (a) Aluminum (b) Water (c) SiGeSn (d) CrFeCoNi.
In the following, we study the performance of three prominent pruning techniques — random, GraNd and CCS — for various pruning factors. ML models for each pruning factor and pruning strategy were trained three times, and the results are reported accordingly. The errors (NRMSE) in electron density predictions for these three pruning approaches are plotted against various pruning factors across all four materials systems in Figure 5. The ML models exhibit minor random fluctuations in their predictions, which is expected due to inherent randomness in the dataset and the model. All three methods perform well at smaller pruning factors across materials, maintaining the accuracy of the original dataset. However, the GraNd method fails catastrophically, with errors rising sharply beyond 40–60% pruning across different materials. Random pruning shows a steady increase in error with pruning factor, preserving the accuracy of the original dataset up to 40–80% and exhibiting rapid error growth beyond that. In contrast, the CCS method outperforms both GraNd and random pruning, maintaining the accuracy of the original dataset up to 80–90% pruning and showing a slower increase in error than random beyond 90%. Remarkably, CCS continues to deliver accurate results at 99% pruning across all materials, with only a modest increase in error compared to the original dataset. The corresponding errors in the total energy of the system, as obtained from post-processing the ML-predicted electron density, are plotted against various pruning factors in Figure 6. For each pruning factor and and strategy, the predicted electron density from one of the three trained ML models was post-processed. All three methods predict the energy nearly as accurately as the original dataset for smaller pruning factors, up to 40%. However, for aluminum all three methods are chemically accurate until 99% pruning, despite a sharp jump in error for the GraNd method beyond 80%. The GraNd method retains chemical accuracy up to 40% pruning for all materials; however, at higher pruning factors, its energy error increases very rapidly for all materials. Random pruning retains chemical accuracy up to 60% pruning for all materials. It maintains chemical accuracy for simple materials, such as aluminum and water, up to 99%, whereas for complex alloys, the limit decreases to 90% for ternary systems and 60% for quaternary systems. The CCS method consistently outperforms the other two approaches, nearly preserving the accuracy of the original dataset up to 90% pruning across all materials. Remarkably, it continues to yield chemically accurate energy predictions even at 99% pruning. Notably, unlike the pure materials considered here — bulk water and aluminum metal — the errors in energies computed for complex alloy systems appear to exhibit random fluctuations with respect to pruning factor for high pruning factors. Thus, for such cases, the errors in the energy, although related with the electron density errors, are not necessarily monotonic with respect to pruning factors. This is consistent with our recent observations for ML based electronic structure prediction of complex quaternary alloys, whereupon the NRMSE in electron density errors alone was found to be not sufficient in completely determining the energy errors [pathrudkar2024HEA]. Nevertheless, we observe that for CCS based pruning, at very high pruning factors, the quaternary systems do achieve chemical accuracy in the energies, while all the other systems achieve significantly higher accuracies. Further details on these pruning results are provided in the supplemental materials (section V.7.)
II.3 Generalizability of the ML models
Figure 7: Model generalizability to unseen configurations: 2D slices of electron density obtained by KS-DFT, ML model trained on the original dataset, 90% CCS, and 99% CCS based pruned datasets. The electron density($\rho$) along the solid line is also compared with KS-DFT, showing remarkable agreement. The energy errors(Ha/atom) are also shown for each case. (Top) Mono-vacancy defect for 32-atoms aluminum systems. (Middle) 216-atoms SiGeSn checkerboard systems (Bottom) Di-vacancy defect for 32-atoms CrFeCoNi systems. The unit of electron density is $\text{Bohr}^{-3}$ in atomic units.
The limited generalization capability of ML models beyond their training data remains a fundamental issue of the field. In particular, the generalization capability of ML models developed using pruned data is yet unexplored. To address this issue, we evaluate the generalization capabilities of the ML models developed here, by assessing their performance on a diverse set of systems that differ significantly from, and were not included in, the training data. These test cases include systems with structural defects such as mono-vacancies, di-vacancies, and handcrafted checkerboard alloy systems featuring species segregation. We find that the generalization capability of the ML models remains unaffected for up to 99% CCS-based pruning.
The ML-predicted densities via CCS (9090% and 9999% pruning) exhibit remarkable agreement with KS-DFT across all of the above test systems, as illustrated in Figure 7. We further compute each system’s energy from ML-predicted electron densities obtained for the original dataset, 90% CCS and 99% CCS based pruned datasets, as illustrated in Figure 7 (see Supplemental Materials, section V.5 for further details). For $99\%$ pruning, we observe chemically accurate predictions for all cases except for the di-vacancies in the quaternary system, where the error is still relatively low ($\mathcal{O}(10^{-3})$ Ha/atom). Additional comparisons between ML-predicted and KS-DFT-computed electron density for various pruning factors, including mono-vacancy defects for the CrFeCoNi system, are presented in the supplemental material(section V.5).
III Discussion
We have explored the hidden redundancy in electronic structure datasets for diverse material systems. To the best of our knowledge, this marks the first exploration of redundancy in electronic structure datasets through systematic pruning approaches. In addition to random pruning, we have employed two ML-based pruning techniques, namely GraNd [paul2021deep] and CCS [zheng2023_CCS]. GraNd based pruning selects a subset of data having the highest importance scores. The CCS approach selects a subset that ensures comprehensive coverage of the dataset based on a probability density-based coverage strategy. These two ML based pruning methods are compared against random grid-point wise pruning — with the goal of revealing the aforementioned redundancy, and helping select the minimal essential dataset without compromising accuracy. Overall, we are able to identify a large fraction of the electronic structure data that is redundant and which can be eliminated from ML training, without compromising the accuracy or the smoothness of the ML predicted fields.
Random grid-point-wise pruning turns out to be an effective strategy, aligning with findings in the computer vision literature. Although its error rate increases steadily, it can outperform several ML-based pruning methods at higher pruning factors  [guo2022deepcore, zheng2023_CCS].
For all four material systems, at lower pruning rates, GraNd captures the redundancy of the dataset effectively, closely matching or outperforming the original dataset. However, at higher pruning rates, GraNd’s performance drops catastrophically. This occurs because, at large pruning factors, most of the easy-to-learn data points are removed, significantly altering the data distribution and leading to poor data coverage.
We found that pruning via CCS consistently outperforms the other two methods, particularly at higher pruning factors. It can systematically identify an order of magnitude smaller subset of the dataset that maintains the performance of the original dataset (i.e., it is practically lossless) for all materials studied here. We have also verified this trend via computing the system’s energy from the ML-predicted electron densities. Furthermore, pruned data sets which are two orders of magnitude smaller than the original (i.e., $99\%$ pruning fraction) also demonstrate surprisingly high accuracy, and result in chemically accurate energy predictions.
We also demonstrate that 9090% and 9999% CCS based pruned datasets retain significant generalization capability of the model, through tests on mono-vacancies, di-vacancies and handcrafted systems that were not included in training. We also compute the system’s energy from the ML-predicted electron densities and demonstrate that the 90% and 99% CCS based pruned datasets can provide chemically accurate predictions.
In addition to identifying the essential subset of the data, pruning can also accelerate ML model training due to the reduced size of the subset. We found that for the materials considered here, ML model training time on the $99\%$ pruned datasets can be lower by $75\%-90\%$, as compared to the training time on the original data (See Section V.6 of the supplemental material). Thus, such pruned datasets can be used to accelerate hyperparameter tuning. In turn, these advancements can be critical in developing ML models that work across chemical space.
IV Methods
The electron density data needed for training, validation, and testing of the ML models was produced via Kohn-Sham Density Functional Theory (KS-DFT) calculations [kohn1965self, Martin_ES] as implemented in the finite-difference based SPARC code [xu2021sparc, xu2020m, ghosh2017sparc]. Norm conserving (ONCV) pseudopotentials [hamann2013optimized, schlipf2015optimization, van2018pseudodojo] and GGA PBE exchange-correlation [perdew1996generalized] were employed. Prior to using the pseudopotentials for electronic structure data generation, tests were performed to verify that lattice parameters and other physical quantities predicted were consistent with the literature and that the energies and forces were sufficiently converged with respect to the mesh fineness. The mesh spacing chosen for the aluminum, water, SiGeSn and CrFeCoNi systems were $0.25$, $0.25$, $0.40$ and $0.20$ Bohr respectively. A tolerance of $10^{-6}$ was chosen for the self-consistent field (SCF) iteration convergence, and the Periodic-Pulay scheme [banerjee2016periodic] was employed for convergence acceleration. The KS-DFT calculations for ML data generation only sampled the gamma point in reciprocal space, as is common in the simulation of large condensed matter systems.
While the aforementioned calculations provided the electron densities for a given atomic configuration, additional techniques were employed to efficiently generate a large sample set of atomic configurations. For the aluminum and SiGeSn alloy systems, Ab Initio Molecular Dynamics (AIMD) simulations were employed, and the electron density data needed for the ML models was extracted from AIMD “snapshots” at regular simulation intervals [pathrudkar2024electronic, zepeda2021deep]. A standard Nosé Hoover thermostat [evans1985nose] with temperatures between $298$ K and $2100$ K was employed for the aluminum systems, while a temperature of $2400$ K was employed for the SiGeSn systems. Fermi-Dirac electronic smearing at a fixed electronic temperature of $631.554$ K was used. As pointed out in our recent contributions, use of thermalized systems to sample the configuration space leads to more generalizable ML models of the electronic structure [pathrudkar2024electronic, pathrudkar2024HEA], due to better sampling of the input features space.
Although AIMD simulations are convenient for data generation, they are computationally costly. In order to reduce the data generation expense for the water and CrFeCoNi alloy systems, atomic configurations were generated through classical Molecular Dynamics (MD) simulations employing machine learning based interatomic potentials [pathrudkar2024HEA]. No accuracy is sacrificed by this method choice since the electron densities corresponding to each configuration are still produced from high-fidelity KS-DFT calculations. At the same time, by running these classical MD simulations at elevated temperatures and over long times, a wide variety of the atomic configurations can be conveniently generated by sampling the resulting MD trajectories. We employed the Materials Graph Library (MatGL) python package [chen2021learning, chen2022universal], which includes the Materials 3-body Graph Network (M3GNet) — a universal machine-learned interatomic potential — for these purposes. The NVT ensemble with Nosé Hoover thermostat was employed for these MD runs as well. The CrFeCoNi alloy data was collected from high-temperature simulations between $2500$ K and $7000$ K, while the water data was obtained from simulations with temperatures between $300$ K and $600$ K.
The timestep for both AIMD and MD simulations was set to $1$ femtosecond. After a brief simulation period to allow the system to thermally equilibrate, atomic configurations (and also the corresponding electron densities, in case of AIMD) were sampled from the simulation trajectories. Additionally, a number of different initial configurations were used to seed these simulations, in order to allow a larger variety in atomic configurations to be sampled.
IV.1 Machine learning map for electron density prediction
ML approaches for predicting electron densities use descriptors that encode the atomic configuration of a local spatial neighborhood as input. Effective descriptors include interatomic distances and angles, as well as the Superposition of Atomic Densities (SAD), both of which have been successful in capturing local atomic environments [chandrasekaran2019solving, pathrudkar2024electronic, pathrudkar2024HEA, li2025image]. The output of these ML models, the electron density, is commonly represented in one of two ways: (1) as sums of atom-centered basis functions [grisafi2018transferable, fabrizio2019electron], which offer computational efficiency but may struggle with complex densities, and (2) as grid-based predictions [zepeda2021deep, fiedler2023predicting], which provide higher accuracy but require significant computational resources due to fine-mesh evaluations.
The present ML model predicts the ground state electron density values at a set of grid points within a computational domain, based on the coordinates and atomic numbers of the atoms. The mapping is computed in two steps: First, atomic neighborhood descriptors are derived for each grid point using the atomic coordinates and species information. Second, a neural network is employed to map these descriptors into the corresponding electron density at each grid point.
As proposed in our recent contribution  [pathrudkar2024HEA], we adopt the position vectors of atoms in body-attached local frames and the atomic species as the descriptors. This formulation ensures that the descriptors are invariant with respect to rotations, translations and species permutations of the system. Moreover, as shown in  [pathrudkar2024HEA] the descriptors maintain approximately the same descriptor-vector size even if the number of species in the system increases, thus ensuring efficient representation of atomic neighborhoods across chemical space. The local frame of reference for calculating the descriptors is obtained using Principal Component Analysis (PCA) of an atomic neighborhood consisting of $M$ atoms. The components of the position vector ($\textbf{r}_{j}$) of the $j$-th atom in the local reference frame are denoted by $(x,y,z)$. The species information is encoded by using the atomic number of the species. The atomic number of the $j$-th atom is denoted as $Z_{j}$. Thus the descriptors for $i$-th grid point ($\mathcal{D}_{i}\in\mathbb{R}^{5M}$) are given as,
$\mathcal{D}_{i}=\bigg\{Z_{j},\|\textbf{r}_{j}\|,\frac{x}{\|\textbf{r}_{j}\|},\frac{y}{\|\textbf{r}_{j}\|},\frac{z}{\|\textbf{r}_{j}\|}\bigg\}_{j=1,\cdots,M}$ (1)
IV.2 Pruning strategies
We briefly recapitulate the pruning strategies used in this work in addition to the random pruning: first, the GraNd score-based pruning [paul2021deep, sorscher2022beyond], followed by a Coverage-centric Coreset Selection (CCS) method [zheng2023_CCS] applied to the GraNd scores.
IV.2.1 Random pruning:
In the present work, a group of grid points from all KS-DFT snap-shots are randomly selected and pruned away. Note that grid points are randomly pruned instead of individual snap-shots.
IV.2.2 GraNd score based pruning:
We consider a supervised regression setting, where the training set is $S=(x_{i},y_{i})_{i=1}^{N}$, drawn i.i.d. from an underlying distribution $P$. Here, $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}^{+}$ denote the descriptors at $i$-th grid point and the electronic charge density at that grid point, respectively. Furthermore, $d$ is the dimension of the atomic-neighborhood descriptors. For a fixed neural network, let the w be the weights and $l$ be the loss of the neural network at a given step. Let $\textbf{w}_{0}$, $\textbf{w}_{1}$, … , $\textbf{w}_{T}$ be iteratively updated weights of the neural network, where, for some sequence of minibatches $S_{0}$, $S_{1}$, …, $_{T-1} S\subseteq S$ of size $M$, we have $\textbf{w}_{t}=\textbf{w}_{t-1}-\eta\sum_{(x,y)\epsilon S_{t-1}}g_{t-1}(x,y)$. Here $g_{t}(x,y)=\nabla_{\textbf{w}_{t}}l(x,y)$, t=1,,Tt=1,...,T, and $\eta$ is the learning rate. The iterations can be perceived as time evolution, with the time derivative of the loss denoted by, $\Delta_{t}((x,y),S_{t})=-\frac{dl}{dt}$. Due to the random initialization of the the neural network, the weight vector at time $t>0$, $\mathbf{w}_{t}$, is considered as a random variable.
[…]
The change in the time derivative of the loss of any example $(x^{*},y^{*})$ resulting from the removal of a training example $(x_{j},y_{j})$ from the dataset $S$ is bounded by the GraNd score associated with the removed example. Formally, for all $(x^{*},y^{*})$, there exists a constant $c$ such that $\big\|\Delta_{t}((x^{*},y^{*});S)-\Delta_{t}((x^{*},y^{*});S_{-j})\big\|\leq c\,\|g_{t}(x_{j},y_{j})\|$. Where $S_{-j}=S\setminus\{(x_{j},y_{j})\}$ and $\Delta_{t}((x^{*},y^{*});S)=-\frac{d\ell}{dt}$. The constant $c$ does not depend on the training examples, $(x,y)$s. Therefore, the contribution of an example $(x,y)$ to change in the loss of any other example $(x^{*},y^{*})$ is bounded by the gradient norm, $\|g_{t}(x_{j},y_{j})\|$, at any time $t$, given the weights $\mathbf{w}_{t}$. In practice, the gradient of the loss, $l$, is computed with respect to the model’s weights $\mathbf{w}_{t}$ at epoch $t$ for each training example. The GraNd score is obtained by computing the expectation over several training runs of $L_{2}$ norm of the gradient. The examples are ranked based on their GraNd scores, with higher scores indicating greater difficulty in learning.
IV.2.3 Coverage-centric Coreset Selection
The purpose of coreset selection is to choose a subset ($\mathcal{S^{\prime}}$) of the data ($\mathcal{S}$) that minimizes the loss on the test set [zheng2023_CCS]. This selection process can be represented through the following optimization problem [sener2018coreset]:
$\min_{\mathcal{S}^{\prime}\subseteq\mathcal{S}:\frac{|\mathcal{S}^{\prime}|}{|\mathcal{S}|}\leq 1-f}\mathbb{E}_{x,y\sim{P}}\left[l(x,y;h_{\mathcal{S}^{\prime}})\right],$ (2)
where $l$ is the loss function, $h_{\mathcal{S}^{\prime}}$ is the model trained with the coreset $\mathcal{S}^{\prime}$, and $f$ is the pruning factor. In other words, $\mathcal{S}^{\prime}$ is a subset of $\mathcal{S}$ which has a size at most (1f)(1-f) of $S$, and which minimizes the expected test loss over the distribution P{P}.
In a classical coverage setting, a set $\mathcal{S^{\prime}}$ is called an r-cover set of $\mathcal{S}$ if a ball of radius $r$ centered at each data point of $\mathcal{S^{\prime}}$ covers the entire $\mathcal{S}$. Zheng et al. [zheng2023_CCS] extended the classical geometric set cover to cover a probability distribution—referred to as the density-based distribution cover. The associated coreset selection method is called the Coverage-centric Coreset Selection (CCS). They also introduced the concept of a partial cover percentage $p$, of the probability distribution $P$, instead of the full cover.
A subset $\mathcal{S}$ of a metric space $(X,d)$ is defined as the p-partial r-cover of a distribution $P$, on the space $X$ if:
$\int_{X}\mathbf{1}_{\cup_{x\in\mathcal{S}}B_{d}(\textbf{x},r)}(\textbf{x})\,d\mu(\textbf{x})=p$
Here $B_{d}(\textbf{x},r)=\{\textbf{x}^{\prime}\in X:d(\textbf{x},\textbf{x}^{\prime})\leq r\}$ denotes a $r$-radius ball centered at x, $p$ quantifies coverage (in percentage), and $\mu$ represents the probability measure associated with the probability distribution $P$. The probability density information over the input space is reflected in the measure $\mu$. The percentage of coverage $p$ increases with the increasing radius $r$ for a given set $\mathcal{S}$ and a distribution $P$. Given a coverage percentage $p$ and a set $\mathcal{S}$, the minimum covering radius $r$ is the smallest value such that $\mathcal{S}$ achieves at least $p$ percentage coverage under distribution $P$. In practice, the distribution $P$ is unknown and so an approximate estimation of the coverage is required. Thus, the area under the p-r curve ($\text{AUC}_{pr}$), is proposed as a metric to assess the coverage of a coreset [zheng2023_CCS]. The quantity $\text{AUC}_{pr}$ defined as the expectation of the minimum distance between the examples that follows the distribution $P$ and the data in $\mathcal{S}$: $\text{AUC}_{pr}(\mathcal{S})=\mathbb{E}_{x\sim P}\left[\min_{x^{\prime}\in\mathcal{S}}d(\textbf{x}^{\prime},\textbf{x})\right]$. The lower the value of the $\text{AUC}_{pr}$, the better the coverage by the coreset.
State-of-the-art methods often tend to eliminate low-importance data points from high-probability density regions, which decreases coverage (increases the $\text{AUC}_{pr}$), and in turn, this can degrade model performance under high pruning rates. In our implementation of CCS, we used the GraNd score as the importance score. However, the coverage obtained by CCS is distinct from that obtained by GraNd based pruning alone. Traditional GraNd removes only low-scoring examples, leading to poor coverage. In contrast, CCS maintains better data coverage in high-probability density regions, reducing $\text{AUC}_{pr}$ and improving performance, even under significant pruning. CCS allocates more sampling budget to high-probability density areas with easy examples, improving coverage of the data distribution. Compared to random sampling, CCS allocates more resources to low-probability density regions, focusing on challenging examples that are critical for training.
CCS steps: Consider a dataset $\mathcal{S}$ = $\{(x_{i},y_{i},s_{i})\}_{i=1}^{n}$ where $s_{i}$ is the importance score for i-th example $(x_{i},y_{i})$. First, CCS selects a pruning rate $\beta$, a tunable hyperparameter representing the percentage of hard examples to remove. Pruning $\mathcal{\beta}\,\%$ of hard examples from $\mathcal{S}$ yields $\mathcal{S}_{o}$. Second, CCS partitions $\mathcal{S}_{o}$ into $k$ strata based on importance (GraNd) scores, using fixed-width score intervals. While each stratum has the same score range, the number of examples per stratum may vary. Third, given the pruning factor $f$, the desired coreset size is determined as m=(1f)nm=(1-f)\,n. This total budget $m$ is distributed evenly across the $k$ strata (another hyperparameter), allocating $m/k$ samples per stratum. If a stratum has fewer examples than its budget, all its examples are included. The final coreset ($\mathcal{S^{\prime}}$) is formed by combining the sampled examples from all strata.
Author contributions
SH, PT and SP worked on developing the framework for pruning and other machine learning aspects. ST and AG worked on the KS-DFT data generation and post-processing calculations. ASB and SG were involved in conceptualization, methodological design, supervision, and securing funding/resources. All authors contributed to writing the manuscript.
Competing Interests
The authors declare no competing interests.
Data Availability
Raw data were generated at Hoffman2 High-Performance Compute Cluster at UCLA’s Institute for Digital Research and Education (IDRE) and National Energy Research Scientific Computing Center (NERSC). Derived data supporting the findings of this study are available from the corresponding author upon request.
Code Availability
Codes supporting the findings of this study are available from the corresponding authors upon reasonable request.
Materials & Correspondence
Correspondence and requests for materials should be addressed to:
Amartya S. Banerjee (email: asbanerjee@ucla.edu), and
Susanta Ghosh (email: susantag@mtu.edu)
Funding
This work was primarily supported by grant DE-SC0023432 funded by the U.S. Department of Energy, Office of Science. This work was partially supported by the National Science Foundation under Grant No. 2442313 and through the UC National Laboratory Fees Research Program of the University of California, Grant Number L25CR9003.
Acknowledgements.
This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231, using NERSC awards BES-ERCAP0033206, BES-ERCAP0025205, BES-ERCAP0025168, and BES-ERCAP0028072. The authors would like to thank UCLA’s Institute for Digital Research and Education (IDRE), the MRI GPU cluster at MTU for making available some of the computing resources used in this work. The authors acknowledge the use of the GPT-5 (OpenAI) model to polish the language and edit grammatical errors in some sections of this manuscript. The authors subsequently inspected, validated and edited the text generated by the AI model, before incorporation.
V Supplemental Materials
V.1 Error Metrics
The normalized root mean square error (NRMSE) of the ML predicted electron density is defined as
$NRMSE=\frac{\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\rho^{i}_{DFT}-\rho^{i}_{ML})^{2}})}{\rho_{DFT}^{max}-\rho_{DFT}^{min}}$ (3)
where N is the number of grid-points for each snapshot. At $i$-th grid point, $\rho^{i}_{DFT}$ represents the electron density calculated using KS-DFT while $\rho^{i}_{ML}$ represents the electron density predicted by the ML models. $\rho_{DFT}^{max}$ and $\rho_{DFT}^{min}$ are the maximum and minimum values of electron density across all grid-points of that snapshot.
Let $u$ is the error field of the ML-predicted electron density, given by $u(x)=\rho_{DFT}(x)-\rho_{ML}(x)$, with $x\in\Omega$. The $H^{1}$ seminorm of the scalar error field $u(x)$ over a domain $\Omega$ is defined as $|u|_{H^{1}(\Omega)}=\left(\int_{\Omega}|\nabla u(x)|^{2}\,dx\right)^{1/2}$. The $H^{1}$ norm is defined as $\|u\|_{H^{1}(\Omega)}=\left(\|u\|_{L^{2}(\Omega)}^{2}+|u|_{H^{1}(\Omega)}^{2}\right)^{1/2}$. These quantities are well defined whenever, $u$ lies in the Sobolev space $H^{1}(\Omega)$.
In the paper, the energy prediction errors are defined as the absolute differences (in Hartree per atom) between KS-DFT energies and the energies obtained by postprocessing the ML-predicted electron density.
V.2 Postprocessing of ML predicted electron density
For application purposes, it is of interest to compute downstream properties from the ML-predicted electron density. Since the Kohn-Sham total ground-state energy is a key property that can be computed from the electron density, it can serve as a useful test for the downstream accuracy of the ML model. We postprocess the density predictions to obtain the corresponding total ground-state energies as follows. The predicted electron density is first rescaled by the total number of electrons; this serves to ensure that the total electronic charge is consistent [pathrudkar2024electronic, briling2021impact, alred2018machine, pathrudkar2022machine]. The scaled electron density $\rho^{\text{scaled}}$ at grid point r is obtained from the predicted electron density $\rho^{\text{ML}}$ via:
$\rho^{\text{scaled}}\left(\textbf{r}\right)=\rho^{\text{ML}}(\textbf{r})\frac{N_{\text{e}}}{\displaystyle\int_{\Omega}\rho^{\text{ML}}(\textbf{r})d\textbf{r}}\,.$ (4)
Here $\Omega$ is the periodic supercell used in the calculations, and $N_{\text{e}}$ is the number of electrons in the system. The rescaled electron density $\rho^{\text{scaled}}$ is then used to set up the Kohn-Sham Hamiltonian and a single diagonalization step is performed by using the same electronic structure calculation framework, as was used in the data generation process [xu2021sparc, xu2020m, ghosh2017sparc]. Thereafter, the energy of the system is computed using the Harris-Foulkes formula [harris1985simplified, foulkes1989tight]. The difference between the energy obtained from the ML-predicted electron density and that obtained from the original DFT density serve as the energy error reported in this work.
V.3 Experimental Details
Dataset details: We use sufficiently large datasets to achieve high accuracy in ML for all four material systems. The KS-DFT dataset for the aluminum system consists of 200 training snapshots (216,000 grid points each), totaling $43.2\times 10^{6}$ grid points, obtained from a 32-atom simulation cell; the test set includes 60 snapshots from six temperatures. The dataset for water includes 112 training snapshots (456,533 grid points each, totaling $51.1\times 10^{6}$ grid points), and 193 test snapshots from four temperatures, obtained from a 32-atom simulation cell. The dataset for SiGeSn system, taken from 45 compositions, consists of 216 training snapshots (totaling $38.6\times 10^{6}$ grid points) and test set includes 90 test snapshots from 45 compositions, obtained from a 64-atom simulation cell. The dataset for CrFeCoNi system, taken from 69 compositions, consists of 140 training snapshots (totaling $40.9\times 10^{6}$ grid points) and test set includes 207 test snapshots from four different temperatures and 69 compositions, also from a 32-atom simulation cell.
Neural Network: We employ a deep neural network with 12 hidden layers, each containing 900 nodes, using GELU activation [hendrycks2016gaussian] in the hidden layers and a ReLU activation in the output layer. A learning rate of 0.001 is used across all cases, with the AdamW optimizer [kingma2014adam, loshchilov2017decoupled].
GraNd score computation: The epoch 10, 30, 10 and 10 are chosen for GraNd score computation for Aluminum, water, SiGeSn and CrFeCoNi respectively. Figure 8 shows the errors on the test dataset for Aluminum, water, SiGeSn, and CrFeCoNi for pruned subsets, for scores computed at different times in training. The Figure 8 shows that the computing scores early in training is sufficient to identify important examples, aligning with the findings of [paul2021deep].
Hyper-parameters for CCS: There are two hyper-parameters for the CCS method. a) the hard cut-off rate ($\beta$) and b) the number of strata ($k$). The optimal $\beta$ generally increases with the pruning rate. The performance of various CCS-based coreset constructed from varying $\beta$ is shown in Figure 10. The error in charge density prediction for various number of strata, $k$ is shown in Figure 9. The error is not very sensitive to the that the number of strata($k$), which is consistent with the findings of Zheng et al. [zheng2023_CCS] for image classification datasets. Hence, in this work, the number of strata($k$) being used is 100 all four material systems for all pruning factors.
Figure 8: Choice of score computation epoch: The error in electron density prediction for (a) Aluminum(20% pruning), (b) Water(40% pruning) (c) SiGeSn(20% pruning) and (d) CrFeCoNi(40% pruning) after computing the GraNd score at different times in training.
Figure 9: Choice of number of strata($k$):The error in electron density prediction with various numbers of strata for Aluminum system for various pruning factors for CCS method.
Figure 10: Choice of hard cut-off rate($\beta$): The error in electron density prediction with varying $\beta$ for (a) Aluminum, (b) water (c) SiGeSn (d) CrFeCoNi for the CCS based pruning. It is observed that the obtained NRMSE with pruned dataset vary with the choice of $\beta$. For higher pruning rate, a higher $\beta$ may be optimal. The points with $\times$ markers failed to beat random pruning.
V.4 Quantifying the coverage of a specific distribution by a subset
The GraNd score and other state-of-the-art importance-based pruning methods often perform poorly at higher pruning rates because of poor data coverage.
We have evaluated the $\text{AUC}_{pr}$ by computing the expected minimum distance from the data in the validation dataset to the data in coreset.
Aluminum
Pruning $\text{AUC}_{pr}$ NRMSE
Rate GraNd CCS GraNd CCS
90% 8.396 8.30 $58\times 10^{-2} 1.1$$15\times 10^{-2}$
99% 8.727 8.584 $2.55\times 10^{-2}$ $1.3\times 10^{-2}$
Water
Pruning $\text{AUC}_{pr}$ NRMSE
Rate GraNd CCS GraNd CCS
90% 23.18 22.93 $3.1\times 10^{-2}$ $1.09\times 10^{-3}$
99% 25.29 24.89 $1.11\times 10^{-1}$ $1.9\times 10^{-3}$
SiGeSn
Pruning $\text{AUC}_{pr}$ NRMSE
Rate GraNd CCS GraNd CCS
90% 76.57 76.30 $2.55\times 10^{-2}$ $8.27\times 10^{-3}$
99% 81.91 80.537 $8.27\times 10^{-2}$ $1.17\times 10^{-2}$
CrFeCoNi
Pruning $\text{AUC}_{pr}$ NRMSE
Rate GraNd CCS GraNd CCS
90% 13.074 13.039 $4.4\times 10^{-2}$ $2.4\times 10^{-3}$
99% 13.416 13.335 $1.27\times 10^{-1}$ $2.88\times 10^{-3}$
Table 1: Comparison of $\text{AUC}_{pr}$ and NRMSE between GraNd and CCS method for different pruning percentages
From the Table 1, we can see that, for 90% and 99% pruning percentage, the CCS has lower $\text{AUC}_{pr}$ than the GraNd method. It confirms that the CCS method based coreset indeed selects a subset that ensures a better data coverage, as evidenced by its lower errors (NRMSE).
V.5 Generalizability
In this section, we provide additional results that demonstrate the generalizability of ML models trained on pruned datasets. The error in electron density prediction for defects in the 32-atoms aluminum system is shown in Figure 11. The generalization error of the GraNd method increases rapidly beyond 8080% pruning, whereas random pruning maintains high accuracy up to 9090% pruning. The CCS method supports a remarkably high pruning factor (9999%) without any significant loss in generalization performance
The errors in electron density prediction for handcrafted checkerboard SiGeSn alloy systems is shown in Figure 12. Checkerboard systems of 216 atoms and 64 atoms were considered. The generalization performance for GraNd worsen rapidly beyond 6060%, whereas random pruning maintains high accuracy up to 8080% pruning factor. However, CCS maintains near-original performance until 9999% pruning.
For CrFeCoNi system with defects, the errors in electron density prediction for defects is shown in Figure 13. The generalization performance for GraNd worsen rapidly beyond 4040%, whereas random pruning maintains high accuracy up to 8080% pruning factors. However, CCS maintains near-original performance until 9999% pruning.
Figure 11: Generalizability tests for Aluminum system: The error in the electron density prediction for various pruning factors for monovacancies for 32-atoms Aluminum system. The shaded region represents the range (maximum to minimum) of three ML models around the mean, shown by a solid line.
Figure 12: Generalizability tests for SiGeSn system: The error in the electron density prediction for various pruning factors for SiGeSn handcrafted checkerboard systems. The shaded region represents the range (maximum to minimum) of three ML models around the mean, shown by a solid line.
Figure 13: Generalizability tests for CrFeCoNi system: The error in the electron density prediction for various pruning factors for (a) mono- (b) di-vacancy defects for 32-atoms CrFeConi system. The shaded region represents the range (maximum to minimum) of three ML models around the mean, shown by a solid line. Figures(c) and (d) shows the magnified view of the Figures(a) and (b)
Figure 14: Error in energy for generalizability tests for CrFeCoNi Monovacancy: The error in the energy from the ML-predicted electron density for the original dataset, 90% and 99% CCS-based dataset.
Figure 15: Model generalizability to unseen CrFeCoNi monovacancy defect configurations: Two-dimensional(2D) slices of electron density obtained by KS-DFT, ML model trained on the original dataset, 90% CCS and 99% CCS based pruned dataset. The electron density along the solid line is also compared with DFT and it shows remarkable agreement with the DFT. The unit of electron density is $\text{e}\cdot\text{Bohr}^{-3}$, where e denotes the electronic charge.
We also compute the total energy of the system from the ML-predicted electron density, via postprocessing, for these generalizability tests. The total energy errors (Hartree/atom) relative to the KS-DFT calculations, are shown in Figure 14, which shows that the even after 99% pruning through CCS, the ML predictions are at the chemically accuracy levels. It demonstrates that the CCS-based pruning, even at two orders of magnitude reduction, retains a remarkably high level of the original model’s generalization capability.
V.6 Timing Comparison
Figure 16: Training time comparison between the ML model trained on the original dataset, 90% CCS and 99% CCS based pruned dataset
Pruned subsets speed up the training process significantly, as shown in Figure 16. Training on $90\%$ CCS-pruned subsets reduces training time (compared to the original dataset) by $70\%$, $84\%$, $74\%$, and $50\%$ for Aluminum, water, SiGeSn, and CrFeCoNi systems respectively, while matching the original models’ predictive performance. $99\%$ CCS-based pruned subsets reduces the training time further, while also suffering nominal increase in NRMSE.
V.7 Additional Results
A complete and unscaled view of the error versus pruning factor for water, SiGeSn, and CrFeCoNi, (corresponding to Figures 5 and 6 of the main text) is provided in Figure 17, Figure 18, and Figure 19 for completeness. It shows that the GraNd method becomes too erroneous beyond 60% pruning for water, SiGeSn and CrFeCoNi. However, the random pruning provides high accuracy up to 90%.
Figure 17: Water system: (a) Error in the predicted electron density for various pruning factors. (b) Error in energy obtained from the predicted electron density for various pruning factors.
Figure 18: SiGeSn system: (a) Error in the predicted electron density for various pruning factors. (b) Error in energy obtained from the predicted electron density for various pruning factors.
Figure 19: CrFeCoNi system: (a) Error in the predicted electron density for various pruning factors. (b) Error in energy obtained from the predicted electron density for various pruning factors.
V.7.1 Isosurface Plots
The 3D electron density fields are shown as iso-surface plots for the four materials in Figures 20, 22, 24, and 26. The overall pattern of electron density iso-surfaces is well captured by the original dataset, 90%, and 99% CCS based pruned datasets.
The corresponding 3D error fields are shown in Figures 21, 23, 25, and 27. Although the error iso-surfaces closely resemble those of the original datasets even after 9090% and 9999% pruning, a slight yet progressive increase in error is observed as pruning levels increase.
Figure 20: Aluminum system at 1500 K. Iso-surfaces of electron density obtained by KS-DFT (a) and ML model for original (unpruned) (b), 9090% CCS (c), and 9999% CCS (d).
Figure 21: Aluminum system at 1500 K. Iso-surfaces of errors in electron density obtained by ML model for original (a), 90% CCS (b), and 99% CCS(c) respectively.
Figure 22: Water system at 600K: Iso-surfaces of electron density (at 0.1 e) obtained by DFT (a) and ML model for original (b), 90% CCS (c), and 99.5% CCS (d). The red and pink spheres represent the Oxygen and Hydrogen atoms.
Figure 23: Water system at 600K: Iso-surfaces of errors in electron density (at 0.0075 e) obtained by ML model for original (unpruned) (a), 90% CCS (b), and 99.5% CCS (c) respectively.
Figure 24: $\text{Si}_{50}\text{Ge}_{12.5}\text{Sn}_{37.5}$ system at 2400K: Iso-surfaces of electron density obtained by DFT (a) and ML model for original (unpruned) (b), 90% CCS (c), and 99.5% CCS (d). The blue, purple, and green spheres represent the Si, Ge, and Sn atoms.
Figure 25: $\text{Si}_{50}\text{Ge}_{12.5}\text{Sn}_{37.5}$ system at 2400K: Iso-surfaces of errors in electron density obtained by ML model for original (a), 90% CCS (b), and 99.5% CCS (c) respectively.
Figure 26: $\text{Cr}_{25}\text{Fe}_{25}\text{Co}_{25}\text{Ni}_{25}$ system at 5000K: Iso-surfaces of electron density obtained by DFT (a) and ML model for original (unpruned) (b), 90% CCS (c), and 99.5% CCS (d). The green, orange, blue, and gray spheres represent the Cr, Fe, Co, and Ni atoms.
Figure 27: $\text{Cr}_{25}\text{Fe}_{25}\text{Co}_{25}\text{Ni}_{25}$ system at 5000K: Iso-surfaces of errors in electron density obtained by ML model for original (a), 90% CCS (b), and 99.5% CCS (c) respectively.